• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng

    Log entry time 16:00:54 on May12,2007

    Entry number 201843

    keyword=day shift summary

    13:55 Moller quad tripped off. We need to either reset it (turn it back on) or retune the beam.
    14:30-end of shift, MCC is trying to tune the beam. Harp scan shows beam spot size too large.

    At the end of shift: raster off, BeO target, waiting for ozone level to drop, waiting for MCC expert to come in;

    Good runs during this shift: